FEL Upgrade Project Weekly Brief –May 13-17, 2002



During a site visit to the manufacturer for the large 180 degree dipole magnets for the recirculations of the FEL Upgrade we witnessed a successful measurement of the flatness of the pole faces: they have been machined to a flatness of <0.0005 inches over the full 3 meter beam path.

Good progress was made this week in surveying stand positions in the FEL vault for the incoming FEL Upgrade hardware.


The project monthly report for April 2002 was completed this week and distributed to the DOE, ONR and AFRL program managers. Cost performance for the IR and UV Upgrade projects was close to the estimate presented at the May 1-3 semiannual review.

On Tuesday, May 14, George Neil and Fred Dylla presented a summary of the FEL development program to the DARPA panel on electric ship systems held at Syntek Systems in Arlington, VA.

The presentation was well received.

On Friday, May 17th, several members of the FEL group and the Accelerator Division participated in a jointly sponsored workshop at the University of Virginia on potential applications of the FEL and future lights sources for analyzing biostructures and nanomaterials.

WBS 4 (Injector):

A third test tube for the W&M implantation system is being polished, various methods of particulate control for the chamber have been implemented. Continued working on PARMELA simulations to meet the e-beam parameters at the match point by finding the proper quadrupoles settings.

The gun stand and lightbox stands positions have been surveyed in preparation for installation. Design work continues on the resistor.

We held a design review for the lightbox for the gun on Wednesday. Drawing signoff expected by the end of next week.



WBS 8 (Instrumentation):

DC Power: The east arc box power supply was connected to the AC lines, and feed for the new upgraded west arc supply is being worked. The old pi bends were run from a single power supply with a 100 amp feed, the new arc supplies are fed from 300 amp each. An additional transformer and panel will be installed for the 12 - 100amp/50 volt power supplies. The transformer/panel will use the 100 amp feed from the old pi bends. The 208 volt power could have been scrounged from the length of the gallery but being able to lock out all supplies from a single location is much safer and saves both labor and materials from using many power panels. A new water manifold is also being installed for the 100 amp'ers

An effort was made at cleaning up noise from the magnet test stand and to identify sources of drift. We have made progress and should be back testing quads by Monday.

Progress is also being made on cabling and VME interfaces. With the installation pace picking up and all that is going on it seams as if 2 days were skipped this week.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

At Magnet Enterprises International (MEI) in Oakland CA. The first eight coils have completed their surface repair. They potted their tenth coil. The white spots have gone away to be replaced by two other puzzling phenomena.

• After being removed from the mold and looking beautifully translucent, the outer surface of the epoxy starts to take on a haze which is theorized as a separation from the glass fibers in zones of the ground wrap between conductors. According to MEI, this happens to coils that are potted with Nadic Methyl Anhydride (NMA) hardener. They usually take care of it with an Acrylic or Silicone sprayed-on coating. (This is why you see painted coils, so the imperfections are hidden) MEI has conducted some tests of treated vs. untreated glass tape with no conclusive results. They are going to try treating a portion of the mold with an alternate, non-silicone based mold release that they can trust and see if the coil in that portion doesn’t have the haze.

- The second phenomenon is a residual of a brown oily residue on the surface of the coil

that is easily removed with xylene. We are in consultation with Jeff Hubrig, our epoxy

consultant and the epoxy manufacturer for advice on both phenomena.

• Upper and lower core’s plates are in final machining. The latest projection is that those parts will be done in five shifts. That puts the assembly to check gap on eight cores the week of May 27. They still have to glue the pole shim material on and paint before final assembly with coils.

• They are going ahead with gluing the Purcell gap shims in place on their practice piece to get the experience and qualify their method.

Injector Dipoles (DU/DV)

• At Master Machine, core pieces are in final machining with first article gap inspection anticipated for the end of next week.

• At Wang NMR, they are working on potting the coils.

Arc Dipoles (GY, GX, GQ)

• Process Equipment Co. (PECo) continued manufacturing the cores and parts of the GQ-GX dipoles. GQ cores are rough machined and going to heat treat. GX cores are going into rough machining

• Wang NMR continues working on the winding fixture for the coils and will set up a winding line separate from the GY coil for these coils.

• AES is drawing up the shim system for the use of these magnets at the highest energy. PECo is aware of the changes and the drawings will be issued in lock step with their need to continue production without rework.

• Wang NMR removed the first GY coil from its form (No small feat). They are reworking their fixture to make the remaining coils easier to wind. Second winding should start next week.

• Bosma Machine complete assembly of the first article GY Core. We inspected it, finding that the gap height was within the .0005 inch (a quarter of a human hair) uniformity required. It is beautiful. After some additional painting, they will ship next week. The second core will be ready for gap inspection the first week of June.


3 inch quad (QX)

• Our decision to re measure all the QX quads to the new CEBAF, bang-bang hysteresis protocol has become more complex. Though the magnet quality is proven, the measurements have been perpetually fogged by drift and noise, Tom Hiatt has asked Kevin Jordan to look at their whole measurement system. The look by an electronic/electrical expert has revealed a legacy of processes and instruments for measuring magnets that are archaic (from the original CEBAF Commissioning days) when compared to present day methods used to run hysteresis loops and set magnets in the CEBAF accelerator. As a result, test protocol is simplified to 17 minutes rather than 45 and will exactly match the magnet setting method. Kevin is tackling the noise and drift as this is being written.

Trim Quad (QT)

• Milhous Control of Virginia continues making the remaining seven magnets.

• No word on the GC corrector because Tom Hiatt is consumed with test stand problems.

Sextupole (SF)

• DULY Research continues to work on completing drawings of the magnet. They are concentrating on designing the splice and cooling water blocks on the coils so that when brazed, the small passage is not blocked and so that the manifolds and connections between the six coils will be compact but serviceable.

Octupole (OT)

• Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

• The drawings for ARC Chambers were signed off.

• The Alignment Group laid out the position of the latest group of stands and a number of the stands in the first arc (They came in this week) were placed in position.

• The recirculation region in front of the Cryomodule is nearing completion.

• The optical chicane region with its chamber is about half done.

• The regions around the wiggler and is in design.

• The only region still not completed or in process is the recirculation dump.

WBS 11 (Optics):

We held informal reviews with optics group staff on the optical cavity vacuum vessels. The few suggestions were incorporated into the design, and the team shifted to producing detailed drawings. We hope to release these next week.

Our vendor for the optical cavity internals requested input on how to fabricate the yaw plate. We are meeting with our vacuum group and experts within the FEL team. We will hold a telephone conference with the vendor next week. Fabrication of the mounts that provides translation for these internal components are in process at the Lab Machine Shop

The same outside vendor is completing the OCMMS fabrication. They report they will complete one assembly this coming week (about 1 week late). Another assembly is delayed due to a slip in delivery of calcium fluoride vacuum viewports.

Mirror mounting tests with sapphire substrates are proceeding. A sapphire mirror mounted using In:Ag braze in A Ni-plated holder was temperature-cycled (warmed to ~ 30 deg. C above ambient, then cooled) with no change in its mounting integrity. A 2" ZnSe mirror's edge was sputter-coated with Ni and will be used for similar mounting tests. Nickel test fixtures that mimic the mounting arrangement on the Upgrade outcoupler holder (but with various gaps) will be used next week to determine the optimum gap required to obtain a uniformly dispensed bead of Ga:In eutectic, another candidate heat conduction medium.

The fabrication of the first article 50kW optical beam dump is nearly complete. A second design, employing "foam metal" (a metal mesh) in the cooling channels to promote turbulent flow and thus increase heat transfer coefficients has been awarded to an outside vendor and is being fabricated.

Our one setup for laser beam profiling is becoming oversubscribed, so we ordered additional hardware and software. The equipment arrived this week and we'll install it for immediate use to relieve the bottleneck. Software that will be used for interpreting laser pulse shapes was installed and tested using simulated data. Initial measurements of the drive laser output after using fiber-grating pulse compression indicate that the pulse has been compressed from a value of ~ 60 ps to a few ps, as expected from reports in the literature.

The chiller used to temperature-stabilize the ultrafast laser system arrived, joining the pump laser. We expect the oscillator next week, which will complete the acquisition of this important piece of hardware. An analysis to determine eye hazards from this laser system was conducted, and appropriate eyewear selected. This was ordered this week. Wiring in UL 1 to provide

safety interlocking for the new laser was also installed this week.

An additional review of the design of the optical component for the THz radiation monitor resulted in a change to return the electro-optical crystal to a position in air. This simplified the design. We hope to complete this design effort in the next two weeks.

Other activities:

Michelle Shinn reviewed the status of large optics and optical subsystems (LRUs) production at the NNSA-led semiannual review of the National Ignition Facility at LLNL. While there, she received an update on the status of magnetorheological finishing of optics, and had some hands-on exposure to the NIF optical mounts.